Suppose we have a binomial experiment in which success is defined to be a particular quality or attribute that interests us. (a) Suppose n = 45 and p = 0.22. (For each answer, enter a number. Use 2 decimal places.) n·p = n·q = Please provide correct answers and show work
Answer #1

solution

given that

n = 45

p = 0.22.

p=0.22=1 - 0.22=0.78

n·p =45 *0.22=9.90

n·q =45*0.78=35.10